OpenStudy (kpop4life123):

please help me :( Ree is 27 inches tall on her first birthday. She grows around 6 inches per year after that. What is the rate of change for the scenario described? (4 points) 6 12 27 33

OpenStudy (anonymous):

D should be the answer ☺

OpenStudy (kpop4life123):

can u help me on one more :)

OpenStudy (anonymous):

sure ☺

OpenStudy (kpop4life123):

A delivery truck has 40,000 miles on its odometer and travels an average of 22,000 miles per year. Which function best models the linear relationship? (4 points) y = 22,000x – 40,000 y = –22,000x + 40,000 y = 22,000x + 40,000 y = 40,000x – 22,000

OpenStudy (anonymous):

i don't know that one hold on let me get help @iGreen @mathmale @mathmath333

OpenStudy (anonymous):

I would pick C.

OpenStudy (anonymous):

Because it is saying the amount of years is multiplied by 22,000 and 40,000 is added to it.

OpenStudy (kpop4life123):

thanks :)

OpenStudy (anonymous):

No problem fam